Solve the system of equations using the addition (elimination) method.
9x + 2y = 2
3x + 5y = 5
Please solve and explain..Thank you

Answer by Prithwis(166)

9x + 2y = 2
3x + 5y = 5 Multiply -3
--------------------------
9x + 2y = 2
-9x-15y=-15
-----------
-13y=-13
y=1
----------
3x+5(1)=5
3x=0
x=o
---------
(x,y)=(0,1)